Gulf States Manufacturing has the following data from year 1 operations, which are to be used for developing year 2 budget estimates:

Sales revenues (14,000 units) $ 1,260,000
Manufacturing costs
Materials $ 225,000
Variable cash costs 307,000
Fixed cash costs 123,000
Depreciation (fixed) 151,000
Marketing and administrative costs
Marketing (variable, cash) 160,000
Marketing depreciation 39,000
Administrative (fixed, cash) 156,000
Administrative depreciation $ 14,000
Total costs $ 1,175,000
Operating profits $ 85,000


All depreciation charges are fixed. Old manufacturing equipment with an annual depreciation charge of $15,750 will be replaced in year 2 with new equipment that will incur an annual depreciation charge of $22,200. Sales volume and prices are expected to increase by 14 percent and 3 percent, respectively. On a per-unit basis, expectations are that materials costs will increase by 12 percent and variable manufacturing costs will decrease by 2 percent. Fixed cash manufacturing costs are expected to decrease by 9 percent.

Variable marketing costs will change with volume. Administrative cash costs are expected to increase by 10 percent. Inventories are kept at zero. Gulf States operates on a cash basis.

1.

value:
25.00 points

Required information

Required:

Prepare a budgeted income statement for year 2. (Do not round intermediate calculations.)

References

eBook & Resources

Financial StatementsLearning Objective: 13-04 Develop production and cost budgets.

Difficulty: 2 MediumLearning Objective: 13-06 Develop budgeted financial statements.

Check my work

2.

value:
25.00 points

Required information

Required:

Estimate the cash from operations expected in year 2. (Do not round intermediate calculations.)
Answer #1
1 Budgeted income statement
$
Sales revenues (Note:1) 1479492
Manufacturing costs:
Materials (Note:2) 287280
Variable cash costs (Note:3) 342980
Fixed cash costs 123000*(100%-9%) 111930
Depreciation (Fixed) (Note:4) 157450
Marketing and administrative costs:
Marketing (Variable,cash) (Note:5) 16279
Marketing depreciation 39000
Administrative (Fixed,cash) 156000*(100%+10%) 171600
Administrative depreciation 14000
Total operating expenses 1140519
Net income 338973
Note:1
Units=14000*(100+14%)=14000*114%=15960
Selling price for year 1=1260000/14000=$ 90 per unit
Selling price for year 2=90*(100%+3%)=90*103%=$ 92.7
Sales revenue=15960*92.7=$ 1479492
Note:2
Material cost per unit for year 1=225000/14000=$ 16.07
Material cost per unit for year 2=16.07*(100%+12%)=$ 18
Units=14000*(100+14%)=14000*114%=15960
Materials cost for year 2=15960*18=$ 287280
Note:3
Variable cost per unit for year 1=307000/14000=$ 21.93
Material cost per unit for year 2=21.93*(100%-2%)=21.93*98%=$ 21.49
Units=14000*(100+14%)=14000*114%=15960
Materials cost for year 2=15960*21.49=$ 342980
Note:4
Increase in depreciation due to replacement of equipment=22200-15750=$ 6450
Depreciation (Fixed) for year 2=151000+6450=$ 157450
Note:5
Variable marketing cost per unit=160000/14000=$ 11.43
Units=14000*(100+14%)=14000*114%=15960
Variable marketing cost for year 2=15960*(100%+2%)=15960*102%=$ 16279
2 Cash from operations expected in year 2:
$
Net income 338973
Add:
Depreciation (Fixed) 157450
Marketing depreciation 39000
Cash from operations 535423
I appreciate your ratings